Grade 12th passAlgebra

There are10 pairs of ear rings. 4 ear rings are selected at random. What is the probability of selecting at least one pair.

sahil suman
9 Years agoGrade 12th pass

1 Answer

DILIP
9 Years ago
There are 20 ear rings in total. We have to select 4 ear rings in random.
In 20 pieces we can select 4 in 20C4 ways.
Total no.of outcomes are 20C4 = 4845.
Our favourable outcomes selecting at least one pair.
one pair can be selected from given 10 pairs in 10C1 ways.
Two pairs can be selected from given 10 pairs in 10C2 ways.
 Total no.of favourable outcomes are 10C1 + 10C2 .= 100 + 45 = 145.
Probability = 145 / 4845 = 29 / 969.